Let x = 2 and y = 3. Find the value of the given capital letter.
N= x^2 +2xy + y^2 + 1
N =
1
Expert's answer
2013-01-23T11:57:07-0500
Let's plug in values of x,y into the formula for N:

N=2^2+2*2*3+3^2+1=4+12+9+1=26

Answer : N=26